OpenStudy (anonymous):

Concentrated nitric acid used in laboratory work is 68% nitric acid by mass in aqueous solution. What should be the molarity of such a sample of the acid if the density of the solution is 1.504 g/ml?

OpenStudy (anonymous):

Concentrated nitric acid used in laboratory work is 68% nitric acid by mass in an aqueous solution. This means that 68 g of nitric acid is dissolved in 100 g of the solution. Molar mass of nitric acid (HNO3) = 1 × 1 + 1 × 14 + 3 × 16 = 63 g mol−1 Then, number of moles of HNO3 Given, Density of solution = 1.504 g mL−1 Volume of 100 g solution = Molarity of solution

OpenStudy (anonymous):

Volume of 100 g solution = 100/1.504 = 66.49 ml = 66.49 * 10^-3L Molarity = 1.079/66.49*10^-3 = 16.23 M